OpenStudy (anonymous):

(help me solve) Sara has some oranges. She sold 40% more than she ate. If she sold 70 oranges, how many did she eat? 25 50 75 100 None of the above.

OpenStudy (whpalmer4):

number eaten = x number sold = 40 % more then x = x + 40% of x = x + 40x/100 number sold = 70 70 = x+40x/100 can you solve for x?

OpenStudy (whpalmer4):

make that "more than"